# Leveraging Content and Acoustic Representations for Efficient Speech Emotion Recognition

### v1

@@ IV-E Results
PASE+ evaluation: We extracted PASE+ features for each downstream dataset, followed by training a two-layer classification network to predict the emotion class of each utterance. The results demonstrate that acoustic arm of the proposed CARE model significantly outperforms the PASE+ baseline across all datasets.
Whisper+RoBERTa Evaluation: In this approach, transcripts are extracted using the Whisper-large-v3 model [19], and subsequently processed by a pre-trained RoBERTa model. The internal representations of the RoBERTa model are linearly combined and a two-layer classification head is trained. It is noted that, with the exception of MELD and CMU-MOSI, the semantic arm of CARE outperforms this baseline for all the other datasets. Further, the Whisper-RoBERTa system contains a large number of parameters ($10$x) compared to the proposed CARE model.
Out-of-Domain Tasks: The performance of CARE is also evaluated on three out-of-domain tasks: RAVDESS song, CaFe [56] and EmoDB [57] speech. The proposed CARE approach significantly improves over all the other models compared. The Whisper-RoBERTa model fails to generalize to these tasks, as the datasets contain song or multi-lingual speech. The results indicate that, in these instances, the acoustic arm of CARE outperforms its semantic counterpart. Despite the fact that neither component of CARE is pre-trained on datasets comprising songs or non-English speech, the model demonstrates strong generalization to these tasks. This success is attributed to the distillation based learning in the proposed model which helps to avoid over-fitting to training domains.
V Discussion
TABLE III: Comparison with other works for downstream datasets. # indicates the models where the downstream dataset is contained in the pre-training dataset. Results in bold indicate the best performing model, while those underlined indicate the second-best model. All numbers are weighted F1-scores computed over $5$ random initializations
Params:$94$M Params:$94$M Params:$94$M Params:$94$M Params:$600$M Params:$8$B Params:$14$B Params:$160$M
IEMOCAP-4 $65.85$ $64.96$ $62.68$ $67.45^{\#}$ $59.43$ $\mathbf{75.78}^{\#}$ $\underline{72.92}^{\#}$ $69.39$
IEMOCAP-6 $51.67$ $50.72$ $46.01$ $54.05^{\#}$ $43.48$ $\mathbf{59.31}^{\#}$ $\underline{58.07}^{\#}$ $55.02$
MELD $45.61$ $45.30$ $41.89$ $47.60^{\#}$ $43.16$ $\mathbf{53.34}$ $\underline{52.59}$ $48.05$
CMU-MOSI $64.09$ $62.53$ $59.67$ $66.45$ $\underline{74.61}$ $\mathbf{78.03}$ $72.80$ $66.74$
DAIC-WOZ $63.15$ $65.87$ $\underline{67.81}$ $61.62$ $64.32$ $62.61$ $64.72$ $\mathbf{68.49}$
RAVDESS-Song $50.45$ $\underline{53.46}$ $38.53$ $48.49$ $11.81$ $50.20$ $51.92$ $\mathbf{60.11}$
CaFE $66.58$ $66.53$ $48.75$ $59.27$ $5.73$ $59.86$ $\underline{69.84}$ $\mathbf{76.98}$
EmoDB $66.54$ $66.94$ $48.87$ $64.37$ $10.2$ $\underline{82.80}$ $82.15$ $\mathbf{83.41}$
V-A Comparison with prior works
A comparative analysis with several recent studies is presented in Table III. These baseline models are categorized into two groups based on the number of parameters used during inference: base models (parameter size $<200$M), and large models ($>500$M), which also include LLM based models. The following observations are made for each category:
V-A1 Base models
We compare HuBERT [15], WavLM [21], data2vec [32] and emotion2vec [42] representations as the baseline models in this category. Among these baseline systems, the emotion2vec is also pre-trained on IEMOCAP and MELD datasets, partially explaining the improved results seen on the downstream tasks on these datasets. While CARE performs similar to emotion2vec on CMU-MOSI, it achieves a significant improvement over all the base-sized models on other datasets. On the average, the proposed CARE achieves a relative improvement of $15.6$% over the best baseline model (HuBERT).
V-A2 Large models
V-A3 LLM based models
Two versions of SALMONN [20] ($8$B and $14$B) [^7] https://huggingface.co/tsinghua-ee/SALMONN are considered as examples of LLM-based models. A direct zero-shot evaluation of the LLM based models achieved poor results. Thus, the same framework used in CARE and other baseline models was followed for the LLM based evaluations as well. The internal representations from all layers ($41$ layers for SALMONN $14$B and $33$ layers for SALMONN $8$B) are aggregated using a convex combination, and the classification head (two-layer feed forward network) is trained for each downstream dataset. Similar to emotion2vec, SALMONN includes IEMOCAP in its pre-training, leading to superior performance on IEMOCAP-4 and IEMOCAP-6 compared to CARE. The larger model size and extensive pre-training data allows SALMONN to outperform CARE by $10\%$ and $34\%$ (relative improvements) on the MELD and CMU-MOSI datasets, respectively. However, on the remaining three tasks, CARE surpasses the SALMONN models, achieving relative improvements of $17\%$ and $24\%$ on the RAVDESS-song and CaFE datasets, respectively.
V-A4 Key takeaways
On average, CARE emerges as the top-performing model across the eight datasets. The SALMONN $13$B model, which has nearly $80$ times more parameters than CARE, is the next best-performing model. Although LLM-based models excel in emotion recognition for some of the in-domain tasks, their performance on the out-of-domain tasks suggests that they may not generalize well on diverse tasks and multi-lingual emotional speech.
@@ V-A4 Key takeaways
TABLE IV: Results on the downstream datasets (weighted F1-score) with modifications in the semantic arm.
@@
IEMOCAP(4-class) CARE-Trans $65.89$
@@
IEMOCAP(6-class) CARE-Trans $51.19$
@@
MELD CARE-Trans $46.57$
@@
CMU-MOSI CARE-Trans $65.97$
@@
DAIC-WOZ CARE-Trans $65.93$
@@
TABLE V: Results on the downstream datasets (weighted F1-score) with different initializations of the acoustic and common arms.
@@
V-B Adapters in the semantic arm
The results in the previous sections used the semantic arm with pre-trained RoBERTa model and convolutional adapters. We make two modifications to this architecture:
@@ V-B Adapters in the semantic arm
    CARE-Trans.: Removing the convolutional adapters while the RoBERTa transformer layers are updated in pre-training.
@@
    CARE-FT: Keeping the convolutional adapters, we update all the parameters (conv. adapters and transformer weights) in the semantic arm.
The results for these two modifications are shown in Table IV. The setup, where the transformers are also trained during pre-training, yields lower performance on the downstream datasets. The results also show that the proposed convolution layer based adaptation strategy offers a good compromise of speech representation learning capability from text-based models and parameter efficiency.
V-C Initialization of Acoustic Arm
As indicated in Section III, the common and the acoustic arm of CARE are initialized with the WavLM-base model weights. We present the results of our method when this initialization is modified to i) random initialization, ii) HuBERT-base [15] or iii) data2vec-base [32] (Table V). The HuBERT-base initialization is seen to outperform the WavLM-base version of CARE for the RAVDESS-song and EmoDB datasets. However, on the average, the WavLM initialization achieves the best performance.
V-D Choice of acoustic targets
We run an ablation experiment where the acoustic arm is trained with targets based on eGeMAPS [9] features extracted from the openSMILE toolkit [25]. The PASE+ targets of the acoustic arm of the CARE model is replaced by the eGeMAPS features. The performance of this model, called CARE (eGeMAPS), is shown in Fig. 2. The performance of the baseline model trained with eGeMAPS as input features is lower than that of the baseline with PASE+ features. Subsequently, the average performance of CARE (eGeMAPS) is lower than of CARE with PASE+ targets.
V-E Continued pre-training of WavLM
To explore the impact of pre-training setup in the proposed CARE, we continued the pre-training of the publicly available WavLM-base model, using the MSP-PODCAST dataset. This was done following the WavLM pre-training procedure, with masked language modeling loss, for an additional $200,000$ steps (similar to the CARE). The results of this experiment are shown in Fig. 3, wherein the performance of the continually pre-trained WavLM model is denoted by WavLM-Cont. Pretrain. The continued pre-training improves the performance of the WavLM-base on some of the downstream tasks like IEMOCAP. However, with the exception of IEMOCAP dataset, the WavLM system’s performance is worse than that of the acoustic arm of the CARE model. As all these models have the same size ($94$M), this ablation highlights the benefits of distillation with the PASE+ representations.
Figure 3: Comparison of the performance when WavLM is continually pre-trained on MSP-PODCAST. The performance of the acoustic arm of CARE is shown for reference. Here, RAVDESS refers to the RAVDESS-Song dataset.
TABLE VI: Results for the MSP-IMPROV dataset. CCC stands for Concordance Correlation Coefficient while V, A, D stand for valence, arousal and dominance respectively.
Method CCC-V CCC-A CCC-D
WavLM-base [16] $0.51$ $0.64$ $0.47$
emotion2vec [42] $0.5$ $0.61$ $0.49$
SALMONN-$8$B [20] $0.53$ $\mathbf{0.67}$ $\underline{0.52}$
SALMONN-$14$B [20] $\underline{0.56}$ $0.65$ $0.51$
CARE $\mathbf{0.57}$ $\underline{0.66}$ $\mathbf{0.53}$
V-F Emotional attribute prediction
The emotion recognition can be posed as a regression problem, where valence, arousal and dominance of a particular utterance is predicted [69]. We use the MSP-IMPROV [70] for this purpose. This is an audio-visual dataset that consists of $12$ actors eliciting a set of sentences in different emotions. The dataset consists of $8438$ utterances with valence, arousal and dominance values (ranging from $1$ to $5$). We first split the dataset in $12$ parts, where each part contains utterances corresponding to $10$ training speakers, while speech from the two other speakers are used for validating and testing the model. The performance is measured as the average over these $12$ parts.
In Eq. 4, $\rho$ is the Pearson’s correlation coefficient between the ground truth and the predicted scores. For training the downstream model, the objective is to increase the CCC between the ground truth and the predicted values for each of the dimensions of valence, arousal and dominance. The results for this dataset along with other baseline models are shown in Table.VI. We note that for this task, the CARE embeddings achieve the best results in terms of the valence and dominance attributes, while the performance on arousal is marginally better for the SALMONN-$8$B model.
TABLE VII: Results on the different datasets for L1111 and L2222 loss
Dataset Loss WF1
IEMOCAP(4-class) L1 $68.54$
L2 $\mathbf{69.39}$
IEMOCAP(6-class) L1 $54.82$
L2 $\mathbf{55.02}$
MELD L1 $47.23$
L2 $\mathbf{48.05}$
CMU-MOSI L1 $66.61$
L2 $\mathbf{66.74}$
DAIC-WOZ L1 $67.53$
L2 $\mathbf{68.49}$
V-G Choice of loss function
We present the results on the different downstream datasets by modifying the loss function to mean absolute loss (L1111) instead of L2222 loss (Table VII). In most cases, the model trained with L2222 loss fares better than the one trained with L1111 loss.
Figure 4: Distribution of cosine similarities for different layer representations for the acoustic and the semantic arm. The top three plots are when two speech signals belonging to different speakers, same emotion and same content are processed by CARE. The bottom three plots refer to the setting with different speakers, same emotion and different content.
V-H Visualization of layers of CARE
In order to interpret the pre-trained model representations learnt by the acoustic and semantic arm of CARE, we probe the representations from the transformer blocks of each arm. We use the English part of the Emotional Speech Dataset (ESD) [71] for this analysis. We form pairs of utterances, where both the utterances of a pair have the same emotional label in all cases and they are derived from two different speakers. A total of $1750$ pairs are considered and the cosine similarities of the pooled representations (for $3$ transformer layers) are shown in Fig.4. The two panels indicate the settings where the speech content in the two utterances are the same (top-panel) or different (bottom-panel). Note that, the representations are obtained from the pre-trained model with unsupervised data, hence we do not expect the model to fully encode emotions. Following are the observations from this visualization:
  1. 1.
    The semantic arm in layer $11$ has slightly more similarity than the acoustic arm for the set of pairs with same content (top-panel).
  2. 2.
@@ V-H Visualization of layers of CARE
    When the spoken content is different, the acoustic arm has significantly higher similarity than the semantic arm, indicating that the acoustic arm is beneficial when the emotion information cannot be reliably predicted from the textual content of the audio.
@@ VI Summary
Key Highlights: In this paper, a pre-training technique for content and acoustic encoding of emotional speech is provided. The proposed architecture, termed CARE, learns an enriched representation of acoustic and semantic information. The acoustic arm uses supervision from low-level descriptors of speech, while the semantic arm is distilled using text representations of the speech transcripts. We also propose a novel adaptation strategy for text-based models in speech representation learning using convolutional neural network layers. The CARE model, with experiments on $8$ downstream tasks, is seen to significantly outperform models of comparable sizes on most of the datasets. Further, the CARE is also observed to generalize better than LLM based models with large parameter sizes. The importance of the different components of the proposed model, along with the different design choices, are established through ablation studies.
Limitations and future scope: The MSP-PODCAST dataset is used for pre-training CARE, which has only $230$ hours of emotional speech data. Training the model on a bigger corpus may improve the performance of the model further. Another limitation of this work, is the relatively lower performance on some in-domain speech datasets compared to LLM-based models, like the CMU-MOSI dataset. In future, we plan to extend the CARE approach to multi-modal speech-text emotion recognition tasks. Further, other masked losses will be explored instead of the L2 based distillation loss utilized in the current work.

### v2

@@ V-A Comparison with Baselines
Baseline $1$: For each downstream dataset, PASE+ features are extracted and a classification network is trained to predict the emotion class of each utterance similar to CARE. The performance of CARE surpasses that of the acoustic supervisory signal by $29.76\%$ (relative) on average across the $8$ datasets. This improvement can be attributed to the larger parameter size of CARE compared to the PASE+ model.
Baseline $2$: In this approach, transcripts are generated for each downstream dataset using the Whisper-large-v3 model [34] and subsequently processed by a pre-trained RoBERTa model. The internal representations from RoBERTa are linearly combined, followed by training a two-layer classification head. Notably, aside from the MELD and CMU-MOSI datasets, CARE outperforms this baseline for all the other datasets, despite the Whisper-RoBERTa system having approximately ten times more parameters. The Whisper-RoBERTa system’s performance on RAVDESS-Song, CaFE, and EmoDB is expected, as these datasets contain identical transcripts spoken or sung with varying emotions, making it difficult for a transcript-based system to distinguish between different emotional utterances.
V-B Importance of the Two Encoders
We present the performance of CARE when we use only one of acoustic and semantic encoders along with the common encoder for the downstream datasets in Table IV. For evaluating the combination of the semantic and common encoders, we use the $768$-dimensional representations from the convolutional feature extractor, the common encoder, and the semantic encoder, excluding outputs from the acoustic encoder. Similarly, the semantic encoder representations are disregarded during the evaluation of the acoustic-common encoder combination. Note that, while CARE has more number of parameters ($160$M) as compared to models like WavLM or emotion2vec, both these combinations have similar number of parameters during inference. While the semantic-common encoder combination has an inference time parameter size of $110$M, the acoustic-common encoder has a total of $94$M parameters during evaluation on each downstream dataset.
Key takeaways: 1) The combination of the acoustic and common encoder representations outperforms the best performing SSL model (HuBERT) by $8.08\%$ (relative) on average for the $8$ datasets (Table II). Given the similar parameter count, this performance suggests an advantage of our pre-training approach. 2) For the three out-of-domain datasets, the acoustic-common combination fares better than its semantic counterpart. This is likely because these datasets contain same sentences with varied emotional expressions, making the acoustic features more effective for distinguishing between utterances. 3) Across all datasets, the combination of both encoders in CARE yields the highest performance, suggesting that while the individual encoder performances are comparable, they capture distinct characteristics of the speech signal.
V-C Modifications in the Semantic Encoder
To evaluate the suitability of our design choices for the semantic encoder, we made three architectural modifications:
  1. 1.
    CARE-No init.: Removing the convolutional adapters, the transformer layers in the semantic encoder are initialized randomly (instead of pre-trained RoBERTa weights).
  2. 2.
    CARE-Trans.: Removing the convolutional adapters while the RoBERTa transformer layers are updated.
  3. 3.
    CARE-FT: Keeping the convolutional adapters, we update all the parameters (conv. adapters and transformer weights) in the semantic encoder.
The results for these modifications are shown in Table V.
Key takeaways: 1) Initializing the transformer weights with RoBERTa is essential for CARE’s performance. Random initialization of the semantic encoder leads to performance drops across all five datasets, suggesting that a randomly initialized semantic encoder struggles to regress to the semantic supervisory signal. 2) Removing convolutional adapters (in CARE-Trans) negatively impacts performance, highlighting the necessity of our convolution-based adaptation technique for aligning speech representations with RoBERTa’s transformer layers. 3) Updating the transformer layers in the semantic encoder decreases performance. Since RoBERTa is pre-trained on text, fine-tuning with speech data degrades its effectiveness.
@@ V-C Modifications in the Semantic Encoder
TABLE V: Results on the downstream datasets (weighted F1-score) with modifications in the semantic encoder.
@@
IEMOCAP(4-class) CARE-No init. $65.71$
CARE-Trans. $65.89$
@@
IEMOCAP(6-class) CARE-No init. $50.96$
CARE-Trans. $51.19$
@@
MELD CARE-No init. $45.16$
CARE-Trans. $46.57$
@@
CMU-MOSI CARE-No init. $62.16$
CARE-Trans. $65.97$
@@
DAIC-WOZ CARE-No init. $64.57$
CARE-Trans. $65.93$
@@
TABLE VI: Results on the downstream datasets (weighted F1-score) with different initializations of the acoustic and common encoders.
@@
V-D Initialization of Acoustic and Common Encoders
As indicated in Section III, the common and the acoustic encoders of CARE are initialized with the WavLM-base model weights. We present the results of our method when this initialization is modified to i) random, ii) HuBERT-base [15] or iii) data2vec-base [58] (Table VI).
Key takeaways: 1) The model’s performance decreases with data2vec initialization, likely due to data2vec’s lower baseline performance compared to HuBERT and WavLM (see Table II). An exception is the CMU-MOSI dataset, where this initialization improves over the WavLM initialized model by $4.21\%$ (relative). 2) The HuBERT-initialized model performs best on the RAVDESS-Song and Emo-DB datasets. Notably, HuBERT outperforms WavLM for these two out-of-domain datasets (Table II). 3) Initialization impacts the acoustic and common encoders less than the semantic encoder, as the latter requires alignment with text representations, making RoBERTa initialization essential.
V-E Choice of Acoustic Targets
We run an experiment where the acoustic encoder is trained with targets based on eGeMAPS [9] features extracted from the openSMILE toolkit [62]. The PASE+ targets of the acoustic encoder of the CARE model is replaced by the eGeMAPS features. The performance of this model, called CARE (eGeMAPS), is shown in Fig. 3.
Key takeaway: The baseline model using eGeMAPS input features performs worse than the baseline with PASE+ features, as expected, since eGeMAPS are handcrafted. Consequently, the average performance of CARE with eGeMAPS is also lower than that of CARE with PASE+ targets.
Figure 4: Comparison of the performance when WavLM is continually pre-trained on MSP-PODCAST. The performance of the combination of acoustic and common encoders of CARE is shown for reference. Here, RAVDESS refers to the RAVDESS-Song dataset.
V-F Continued Pre-training of WavLM
A key question in assessing the effectiveness of our proposed pre-training methodology is the influence of the MSP-PODCAST pre-training dataset. Since all self-supervised learning (SSL) models are trained on neutral data, their ability to accurately discern emotions from speech signals is typically limited. The emotion recognition performance of these SSL models when pre-trained on emotion datasets thus becomes crucial. To explore the impact of pre-training setup in the proposed CARE, we continued the pre-training of the publicly available WavLM-base model, using the MSP-PODCAST dataset. This was done following the WavLM pre-training procedure, with masked language modeling loss, for an additional $200,000$ steps (similar to the CARE). The results of this experiment are shown in Fig. 4, wherein the performance of the continually pre-trained WavLM model is denoted by WavLM-Cont. Pretrain. The performance of the combination of the common and acoustic encoders is also shown for comparison.
Key takeaway: Continued pre-training improves WavLM-base performance on certain downstream tasks, like IEMOCAP. However, except for IEMOCAP, WavLM performs worse than CARE’s acoustic-common encoder combination. As all the models in Fig. 4 have the same size ($94$M) during inference, this experiment highlights the benefits of our proposed distillation based pre-training.
Figure 5: Distribution of cosine similarities for different layer representations for the acoustic and the semantic encoders. The top three plots are when two speech signals belonging to different speakers, same emotion and same content are processed by CARE. The bottom three plots refer to the setting with different speakers, same emotion and different content.
V-G Visualization of Layers of CARE
In order to interpret the pre-trained model representations learnt by the acoustic and semantic encoder of CARE, we probe the representations from each encoder. We use the English part of the Emotional Speech Dataset (ESD) [63] for this analysis. We form pairs of utterances, where both the utterances of a pair have the same emotional label in all cases and they are derived from two different speakers. A total of $1750$ pairs are considered and the cosine similarities of the pooled representations (for $3$ transformer layers) are shown in Fig.5. The two panels indicate the settings where the speech content in the two utterances are the same (top-panel) or different (bottom-panel). Note that, the representations are obtained from the pre-trained model with unsupervised data, hence we do not expect the model to fully encode emotions. Following are the observations from this visualization:
@@ V-G Visualization of Layers of CARE
    The semantic encoder in layer $11$ has slightly more similarity than the acoustic encoder for the set of pairs with same content (top-panel).
@@
    When the spoken content is different, the acoustic encoder has higher similarity than the semantic encoder, indicating that the acoustic encoder is beneficial when the emotion information cannot be reliably predicted from the textual content of the audio.
@@ VI Summary
Key Highlights: In this paper, a pre-training technique for content and acoustic encoding of emotional speech is provided. The proposed architecture, termed CARE, learns an enriched representation of acoustic and semantic information. The acoustic encoder uses supervision from low-level descriptors of speech, while the semantic encoder is distilled using text representations of the speech transcripts. We also propose an adaptation strategy for text-based models in speech representation learning using convolutional neural network layers. The CARE model, with experiments on $8$ downstream tasks, is seen to outperform models of comparable sizes on most of the datasets. Further, the CARE is also observed to generalize better than LLM based models with large parameter sizes. The importance of the different components of the proposed model, along with the different design choices, are established through ablation studies.
Limitations and future scope: The MSP-PODCAST dataset is used for pre-training CARE, which has only $230$ hours of emotional speech data. Training the model on a bigger corpus may improve the performance of the model further. Another limitation of this work, is the relatively lower performance on some in-domain speech datasets compared to LLM-based models, like the CMU-MOSI dataset. In future, we plan to extend the CARE approach to multi-modal speech-text emotion recognition tasks.